Cal-Z Shoes produces shoes in Leon, Guanajuato, Mexico. They operate out of a 3,000 M2 facility, of which 2,000 M2 is the main production area. They have two main product lines. Their own Big Apple shoe brand, for institutional and personal use, are sold to retail outlets in Mexico and the USA and account for 40% of the company’s total production. The second product line is casual shoes for a larger shoe company that sell under that company’s brand name.

Highlights of cost information from the 2019 budget (in Mexican Pesos) includes:

Expense

Total in pesos

Expense

Total in pesos

Shop floor wages

$ 7,000,000

Variable overhead

$12,500,000

Leather

$25,000,000

Fixed overhead

$10,000,000

Other materials

$ 7,500,000

Selling & Admin – Variable

$10,000,000

Selling & Admin – Fixed

$15,000,000

  • Total budgeted production and sales are based on 2,000 pairs of shoes per day for a total of 250 days.
  • Shop floor workers’ wages: 80% are direct labour; the remainder are indirect costs relating to setup and batch preparation. There are 200 production workers.
  • Raw materials: Leather (note 40% is scrapped due to cutting and quality).
  • Average selling price is $200 pesos.
  • $1.00 Mexican Pesos is worth approximately $0.068 Canadian.

What is the margin of safety (in units)?

What is budgeted profit / loss for 2019 (in pesos)?

What is the average hourly wage (in CDN $). Assume 45 hours per week, 50 weeks.

Assume retailers take a 50% margin on sales. What is the average retail price in CDN $?

Assume the USA applies a tariff of 25% on imported shoes from Mexico. What is the average cost of this tariff per pair of shoes?

Assume overhead costs are absorbed on a per-unit basis. What is the budgeted overhead rate?

Assume the factory has unused capacity. A shoe retailer in Leon has approached you with an offer to buy 50,000 pairs of shoes at a price of $175 pesos. Should you take the offer? Why?
Answer #1

MARGINAL COSTING

CALCULATION OF MARGIN OF SAFETY ( IN UNITS) :

Margin of safety = total sales - BEP ( break even point ) sales

  Total sales = 2000 pair of shoes / day * 250 days

= 500,000 pair of shoes

Break even point ( in units) = fixed cost / contribution per unit

(NOTE: break even point is where there is neither profit nor loss )

Fixed cost = fixed overhead + selling and adm oh

= 10,000,000 + 15,000,000

= 25,000,000

contribution per unit = selling price per unit - variable cost per unit selling price per unit (given) = 200$

variable cost per unit : ( material + labour + overheads ) / 500,000

leather (25,000,000) + material (7,500,000) + wages (7,000,000)

+ overhead (12,500,000) + selling & adm (10,000,000)

= 62,000,000 / 500,000

= 124$

contribution per unit = 200 - 124 = 76$

BEP ( in units ) = 25,000,000 / 76

= 328,947 units

MARGIN OF SAFETY = 500,000 - 328,947

   MOS = 171,053

CALCULATION OF BUDGETED PROFIT / LOSS :

sales (200 * 500,000) = 100,000,000

( - ) variable cost =( 62,000,000)

contribution = 38,000,000

(-) fixed cost = ( 25,000,000)

PROFIT = 13,000,000

CALCULATION OF HOURLY WAGE RATE :

= wages (in CND $ ) / number of hours

= 476,000 / 450,000

= 1.058$ (CND ) per hour per labour

WORKING NOTE :

wages in CND $ = 7,000,000*.068

= 476,000

number of hours = 45hrs per week * 50 weeks * 200 workers

= 450,000 hours

CALCULATION OF AVG RETAIL PRICE :

retailers sell at 200$ with margin of 50% on 200$

200$*50% = 100 $

RETAIL PRICE = 200$ - 100$

= 100$ * .068

AVG RETAIL PRICE = 6.8 CND $

CALCULATION OF TARIFF PER PAIR OF SHOE :

   25% on imported shoes = 200$ * 25%

= 50$ per pair of shoe

CALCULATION OF BUDGETED OVERHEAD RATE :

   = ( variable overhead + fixed overhead ) / number of shoes

VOH ( including selling & admin) = 22,500,000

FOH (including selling & admin) = 25,000,000

= 47,500,000/ 500,000

=  95 $ per shoe

DECISION MAKING :

Shoes must not be sold to the retailer in Leon .

normal price = 200$

price to retailer = 175$

LOSS = 25$

So , the 50,000 shoes must not be sold to the retailer .